OpenStudy (anonymous):

if a + b + c = -1a+b+c=−1 and x + y + z = 3x+y+z=3, what is -7a + 2z + 2y + 2x - 7c - 7b−7a+2z+2y+2x−7c−7b?

OpenStudy (dayakar):

a+b+c= -1--------(I) x+y+z= 3---------(II) given

OpenStudy (dayakar):

write all the in order

OpenStudy (dayakar):

=-7a-7a-7b-7b-7c-7c+2x+2x+2y+2y+2z+2z

OpenStudy (dayakar):

add all like terms

OpenStudy (dayakar):

=-14a-14b-14c+4x+4y+4z

OpenStudy (dayakar):

r u following

OpenStudy (dayakar):

= -14(a+b+c) + 4(x+y+z) [taking -14,and 4 as common]

OpenStudy (dayakar):

=-14*(-1) + 4* 3 {from (I) and (II) which is given } =14 +12 =26

OpenStudy (dayakar):

r u clear